Please help!!

A random sample of 15 size AA batteries for toys yield a mean of 3.18 hours with standard deviation, 0.52 hours.


(a) Find the critical value, t*, for a 99% CI. t* =

(b) Find the margin of error for a 99% CI
Answer #1

Solution(a)
Given no. of sample = 15 and degree of freedom = (15-1) = 14
and alpha = 0.01 and alpha/2 = 0.005
from t table we found t= 2.9768
Solution(b)
margin of error = t*SD/sqrt(n) = 2.9768*0.52/sqrt(15) = 0.3996